Suppose that .25 cubic inch of paint is applied to a circular piece of plastic with a diameter of 20 inches. Estimate the thickness of the paint.
Answer by Alan3354(69443)

Suppose that .25 cubic inch of paint is applied to a circular piece of plastic with a diameter of 20 inches. Estimate the thickness of the paint.
------------
r = 10 inches
Area = pi*r^2 = 314 sq inches
T = 0.25/314 =~ 0.000796 inches